40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Sep 8 17:48:10 2017, Kira, DailyProgress, TempCtrl, Heater Circuit noise spectrum.pdfIMG_20170908_172648.jpg
    Reply  Wed Sep 13 14:17:20 2017, awade, DailyProgress, TempCtrl, Re: Heater Circuit noise 
       Reply  Thu Sep 14 17:56:47 2017, Kira, DailyProgress, TempCtrl, Re: Heater Circuit noise 
Message ID: 1916     Entry time: Wed Sep 13 14:17:20 2017     In reply to: 1904     Reply to this: 1917
Author: awade 
Type: DailyProgress 
Category: TempCtrl 
Subject: Re: Heater Circuit noise 

Edit (awade Wed Sep 13 17:40:35 2017): Gautam may also have suggestions on how to do this better

Great!

We might also want to characterizing the input DAC noise into your circuit.  I think Yinzi and I did this a while ago, I think she made a post to the elog.  Maybe you'd like to re do this.  It basically amounts to outputting a slow (10- 100 mHz) sine wave to tickly the bits from the DAC and measuring the voltage noise (directly from the DAC) with the SR785.  We need to understand this noise in order to understand how much noise the heater driver is adding and how much is due to our DAC.

You might also want to compute the total rms noise in the band of interest by integrating the PSD (ASD?) down from high frequency down to low. From this you can work out the uncertainty/noise in the heat actually delivered. 

---

As a bonus, it can be a good idea to zip or  tar your data and the scripts used to plot it and attach to posts. Its a little more work at the time but it means you can easily work out what is going on in 6 months time.

Craig and I can also give you access to the ctn_labdata git to commit it to our data repository. 

 

 

Quote:

Kevin and I took measurements of the heater circuit noise using a spectrum analyzer. The heater was attached to it by both ends using BNC cables clipped to the ends of the heater. We used 600mV for the DAC voltage and +24V/-24V for the input. The temperature of the metal block changed from 34.1C to 34.7C as we took the measurement, which took about 13 minutes. We took an RMS average of 25 measurements in total to get the spectrum, shown in the first attachment. The setup is shown in the second attachment.

We first took a range of 0-200Hz for our measurement, but it was not accurate enough to see the dependence of the ASD and the frequency as the lower frequencies didn't have enough data points to accurately plot it. We then took 0-50Hz and got the result shown.

We also mounted the heater circuit box to the rack.

 

ELOG V3.1.3-